OpenStudy (anonymous):

Simplify: x^2-10x+9 ----------- x^2+2x-3

OpenStudy (hoblos):

x^2-10x+9 = (x-1)(x-9) x^2+2x-3 = (x-1)(x+3)

OpenStudy (hoblos):

x^2-10x+9 x-9 ----------- = --------- x^2+2x-3 x+3

OpenStudy (ash2326):

We have \[\frac{x^2-10x+9}{x^2+2x-3}\] Let's find the factors of \(x^2-10x+9\) we have to find factor of 9, such that the sum is -10 -1 and -9 satisfy this so \[x^2-10x+9=x^2-9x-x+9\] we have now \[x(x-9)-1(x-9)=(x-1)(x-9)\] so \[x^2-10x-9=(x-1)(x-9)\] Now let's find the factor of \(x^2+2x-3=0\) we have to find the factors 0f -3 such that their sum is 2 3 and -1 satisfy this we get \[x^2+2x-3=x^2+3x-x-3=>x(x+3)-1(x+3)\] we get \[x^2+2x-3=(x-1)(x+3)\] We have \[\frac{x^2-10x+9}{x^2+2x-3}=\frac{\cancel{(x-1)}(x-9)}{\cancel{(x-1)}(x+3)}\] we get \[\frac{x-9}{x+3}\]